# WoW: Towards a World-omniscient World-model Through Embodied Interaction

### v1

@@ 4.1.1 Pretrain Data Preparation
We construct our training dataset through a multi-stage pipeline designed to ensure both quality and diversity. The process consists of four sequential stages: Collection, Filtering, Refinement, and Rebalancing. Unlike simply enlarging the dataset with indiscriminate samples, our approach emphasizes that data quality plays a decisive role in model performance, and carefully curated data prove more effective than raw scale (Collaboration et al., 2023; Khazatsky et al., 2024; Radford et al., 2021; Northcutt et al., 2021; Zhang et al., 2017; Luo et al., 2024).
@@ 4.1.2 Diffusion-Based Video Generation
Auxiliary Perception. To strengthen initial state understanding, we inject features from a self-supervised visual representation model (DINOv2 (Oquab et al., 2023)) into intermediate layers of the DiT. These semantically grounded features improve pixel-level reasoning about object boundaries and spatial relationships, compensating for potential weaknesses in latent representations learned solely through noisy reconstruction objectives.
@@ 4.2 Solver-Critic Video Generation Agents
Building upon the conceptual framework of WoW, this section details the Experience Reflection (Perception) stage—a key component of our closed-loop mechanism. We introduce a solver-critic paradigm that enhances the physical consistency and realism of generated outputs. Unlike Simple ability of previous work (Soni et al., 2024; Chi et al., 2025a), WoW has a comprehensive agent system, driven by a Refiner (Section 4.2.2), which, guided by a Dynamic Critic Model Team (Section 4.2.3), to iteratively improve the generated content. This dynamic workflow forms a Closed-Loop Generative Workflow (Section 4.2.4), ensuring that the model’s output is continuously refined and verified.
@@ 4.2.1 Framework Overview
This architecture transforms the model from a passive generator into an active problem-solver. Our solver-critic framework is built upon three core components: the Refiner Agent, which optimizes the input and generate video output; the Dynamic Critic Model, which evaluates the generated output; and the integrated closed-loop Workflow. Furthermore, we discuss how this architecture aligns with the Prover-Verifier paradigm (Kirchner et al., 2024), showcasing its potential to endow the generative process with a new level of cognitive depth.
@@ 4.2.2 Refiner Agent in World Model
The quality of a generative model’s output is highly dependent on its input prompt. For video generation in specialized fields like robotics, prompts must capture subtle physical details to produce plausible outcomes. However, manually crafting such high-quality prompts is a time-consuming and arduous process of trial and error. While the emerging field of Automatic Prompt Engineering offers systematic optimization methods for language tasks (Khan et al., 2025; Agrawal et al., 2025), these general approaches are not directly tailored to the unique demands of physically-grounded video synthesis.
@@ 4.3 Flow-Mask Inverse Dynamics Model
The proposed Flow-Mask Inverse Dynamics Model (FM-IDM) is a video-to-action model that maps predicted video frames to real-world robot execution transitions. Instead of relying on model-specific features (Liao et al., 2025; Chi et al., 2025b; Hu et al., 2025b), we adopt a pixel-level decoding approach, trading real-time performance for greater generality and accuracy (Ko et al., 2023; Tan et al., 2025). Designed as a plug-and-play module, our model is compatible with any visual generative world model, enabling system-level evaluation and facilitating reward extraction via embodied interaction.
@@ 5.1 Core Evaluation Dimensions
Embodied agents must execute long-horizon tasks (Chen et al., 2024b; Sermanet et al., 2023; Li et al., 2024; Cheng et al., 2024a; Yang et al., 2025b). Therefore, we assess a model’s planning ability by challenging it to generate coherent video sequences for complex instructions. This requires implicitly understanding task decomposition (Chi et al., 2025a; Tian et al., 2025) into key sub-goals and respecting their causal dependencies. Hence, we collect 25 samples to fill in this long-term planning task, and transform the text instruction into a suitable description for the world model to plan. In order to evaluate the planning ability in the world model, we refer to the metric from RoboBench. As illustrated in the center-left of Figure 9, first, we extract the key steps from the predicted video using Gemini-2.5-flash (Comanici et al., 2025), for instance, 1-grasp(green block), 2-pick up(green block), 3-place(green block, yellow block), and so on, then transform them into a Directed Acyclic Graph(DAG) for further key action extraction and comparison with groundtruth DAG. The detailed metric will be extended in Section 5.3.
@@ 5.4 Overall Benchmark Score
Overall score.
For each model $i$ and metric $m$, we map the raw value $x_{i,m}$ to a common desirability score $s_{i,m}\in(0,100)$, then aggregate by groupwise geometric means.
Normalization. Let $F_{m}$ be the empirical CDF (computed on a reference set).
(ECDF, higher is better) $s_{i,m}=100\,F_{m}(x_{i,m}),$ (9)
(ECDF, lower is better) $s_{i,m}=100\bigl(1-F_{m}(x_{i,m})\bigr).$ (10)
For PSNR we use a piecewise power mapping with anchors $L=p_{10},\ U=p_{90}$:
$s_{i,\text{PSNR}}=\begin{cases}0,&x_{i,\text{PSNR}}\leq L,\\ 100\left(\dfrac{x_{i,\text{PSNR}}-L}{U-L}\right)^{0.7},&L<x_{i,\text{PSNR}}<U,\\ 100,&x_{i,\text{PSNR}}\geq U.\end{cases}$ (11)
For error/distance metrics (e.g., ATE, RPE; lower is better) we apply a robust $z$-score using median $\tilde{\mu}$ and $\tilde{\sigma}=1.4826\,\mathrm{MAD}$, followed by a logistic:
$z_{i,m}=\frac{\tilde{\mu}_{m}-x_{i,m}}{\tilde{\sigma}_{m}},\qquad s_{i,m}=100\,\sigma\!\left(\gamma z_{i,m}\right),\quad\gamma=\frac{\mathrm{logit}(0.9)}{2}\approx 1.10.$ (12)
For bounded scales (e.g., execution, physical_law, dag on $0\text{--}5$):
$s_{i,m}=100\cdot\frac{x_{i,m}}{5}.$ (13)
All $s_{i,m}$ are finally clipped to $[2,98]$ to avoid boundary effects.
Correlation-aware intra-group weights. Metrics are organized into groups $g$ (e.g., quality, instruction, physical, planning). Using the correlation matrix $R$ of $s_{i,m}$} within group $g$, we shrink redundant metrics via
$p_{m}=\sum_{j\neq m}\!\lvert R_{mj}\rvert,\qquad\tilde{w}_{m|g}=\frac{1}{1+p_{m}},\qquad w_{m|g}=\frac{\tilde{w}_{m|g}}{\sum_{k\in g}\tilde{w}_{k|g}},$ (14)
with a uniform fallback if $R$ is degenerate.
Aggregation (geometric means). The group score for model $i$ is
$G_{i,g}=100\prod_{m\in g}\left(\frac{s_{i,m}}{100}\right)^{w_{m|g}},$ (15)
and the overall score is
$O_{i}=100\prod_{g}\left(\frac{G_{i,g}}{100}\right)^{W_{g}},\qquad\sum_{g}W_{g}=1.$ (16)
We handle missing values by re-normalizing $w_{m|g}$ over available metrics for each sample, and re-normalizing $W_{g}$ over available groups. Computations are performed in log-space with a small $\varepsilon$ to ensure numerical stability.
@@ 6.2 Model Comparsion Experiment
Cogvideo cogvideo 3.29 1.52 1.73 1.30 7.84 18.93 5.91 44.13 2.32 10.34
Cosmos-Predict1 cosmos1 2.84 2.60 2.41 2.49 10.34 12.39 45.29 16.85 7.47 16.30
Wan2.1 wan 3.49 1.79 2.30 1.62 9.21 12.92 12.63 5.96 5.6 8.59
Cosmos-Predict2 cosmos2 3.18 2.33 2.31 1.62 9.21 34.92 16.52 28.76 6.67 18.24
@@
cosmos1 + Agent cosmos1 3.34 48.26 5.05 8.23 9.05
[…]
@@ 6.4.1 Visual Evaluation
We qualitatively evaluate the generalization capabilities of our World Foundational Model across three critical axes: robotic embodiment, task repertoire, and other variations. Cross-embodiment generalization is demonstrated in Figure 14. Proposed world foundational model successfully follows instructions across a diverse set of embodied hardware platforms without any fine-tuning. The model is validated on a wide range of robotic platforms, including the UR5 and Franka industrial manipulators, the Franka Sim simulation, the Agilix parallel-kinematic arm, and the complex Tiangong dexterous hand. This result substantiates our model’s ability to learn an embodiment-agnostic representation of physical interactions, decoupling the task goal from the specific kinematics and dynamics of the embodied platform. Second, Figure 15 illustrates the model’s task generalization, showcasing its proficiency across a repertoire of 15 distinct manipulation skills that range from simple primitives like pull, push, and move, to complex, contact-rich behaviors such as press button, unstack, and tie. The model’s proficiency across this action space highlights its capacity to learn a compositional skill representation rather than merely memorizing individual task solutions. Figure 16 elucidates the model’s remarkable robustness to profound domain and attribute variations. It maintains successful task execution across drastically different visual styles, including photorealistic, pencil sketch, and oil painting. Furthermore, it adeptly handles varied object properties, manipulating both rigid bodies and fluids, from extremely small to regular-sized objects. The model also exhibits invariance to initial state configurations, consistently picking a red pen from a cup regardless of its spatial pose. Collectively, these visualizations provide compelling evidence that the proposed model acquires a truly abstract and foundational understanding of the world, reasoning about underlying physical principles rather than overfitting to superficial contextual cues.
@@ 6.4.2 Quantity Evaluation
Building upon the preceding qualitative analysis, we conduct a rigorous quantitative evaluation across the benchmark’s diverse material properties and physical phenomena. As detailed in Table 4, we assess four models using a suite of five metrics including FVD, WorldScoreBench, PhysGen, DreamSim, and execution quality score. The results unequivocally demonstrate the superiority of Cosmos and WAN over the other models. This performance gap is particularly pronounced in challenging scenarios involving complex dynamics. For instance, Cosmos achieves the highest PhysGen scores, outperforming the next-best model by over 5% on tasks with deformable objects, which suggests a more accurate internal representation of non-rigid body mechanics. In parallel, world foundational model based on Cosmos consistently records the lowest FVD scores in fluid and optical physics simulations, indicating a state-of-the-art capability for rendering visually realistic and temporally coherent complex scenes.

### v2

@@ 4.1.1 Pretrain Data Preparation
We construct our training dataset through a multi-stage pipeline designed to ensure both quality and diversity. The process consists of four sequential stages: Collection, Filtering, Refinement, and Rebalancing. Unlike simply enlarging the dataset with indiscriminate samples, our approach emphasizes that data quality plays a decisive role in model performance, and carefully curated data prove more effective than raw scale (Collaboration et al., 2023; Khazatsky et al., 2024; Radford et al., 2021; Northcutt et al., 2021; Zhang et al., 2017; Luo et al., 2024; An et al., 2024).
@@ 4.1.2 Diffusion-Based Video Generation
Auxiliary Perception. To strengthen initial state understandingYu et al. (2025), we inject features from a self-supervised visual representation model (DINOv2 (Oquab et al., 2023)) into intermediate layers of the DiT. These semantically grounded features improve pixel-level reasoning about object boundaries and spatial relationships, compensating for potential weaknesses in latent representations learned solely through noisy reconstruction objectives.
@@ 4.2 Solver-Critic Video Generation Agents
Building upon the conceptual framework of WoW, this section details the Experience Reflection (Perception) stage—a key component of our closed-loop mechanism. We introduce a solver-critic paradigm that enhances the physical consistency and realism of generated outputs. Unlike Simple ability of previous work (Soni et al., 2024; Chi et al., 2024), WoW has a comprehensive agent system, driven by a Refiner (Section 4.2.2), which, guided by a Dynamic Critic Model Team (Section 4.2.3), to iteratively improve the generated content. This dynamic workflow forms a Closed-Loop Generative Workflow (Section 4.2.4), ensuring that the model’s output is continuously refined and verified.
@@ 4.2.1 Framework Overview
This architecture transforms the model from a passive generator into an active problem-solver. Our solver-critic framework is built upon three core components: the Refiner Agent, which optimizes the input and generates video output; the Dynamic Critic Model, which evaluates the generated output; and the integrated closed-loop Workflow. Furthermore, we discuss how this architecture aligns with the Prover-Verifier paradigm (Kirchner et al., 2024), showcasing its potential to endow the generative process with a new level of cognitive depth.
@@ 4.2.2 Refiner Agent in World Model
The quality of a generative model’s output is highly dependent on its input prompt. For video generation in specialized fields like robotics, prompts must capture subtle physical details to produce plausible outcomes. However, manually crafting such high-quality prompts is a time-consuming and arduous process of trial and error. While the emerging field of Automatic Prompt Engineering offers systematic optimization methods for language and generation tasks (Khan et al., 2025; Agrawal et al., 2025; An et al., 2025), these general approaches are not directly tailored to the unique demands of physically-grounded video synthesis.
@@ 4.3 Flow-Mask Inverse Dynamics Model
The proposed Flow-Mask Inverse Dynamics Model (FM-IDM) is a video-to-action model that maps predicted video frames to real-world robot execution transitions. Instead of relying on model-specific features (Liao et al., 2025; Chi et al., 2025; Hu et al., 2025b), we adopt a pixel-level decoding approach, trading real-time performance for greater generality and accuracy (Ko et al., 2023; Tan et al., 2025). Designed as a plug-and-play module, our model is compatible with any visual generative world model, enabling system-level evaluation and facilitating reward extraction via embodied interaction.
@@ 5.1 Core Evaluation Dimensions
Embodied agents must execute long-horizon tasks (Chen et al., 2024b; Sermanet et al., 2023; Li et al., 2024; Cheng et al., 2024a; Yang et al., 2025b). Therefore, we assess a model’s planning ability by challenging it to generate coherent video sequences for complex instructions. This requires implicitly understanding task decomposition (Chi et al., 2024; Tian et al., 2025) into key sub-goals and respecting their causal dependencies. Hence, we collect 25 samples to fill in this long-term planning task, and transform the text instruction into a suitable description for the world model to plan. In order to evaluate the planning ability in the world model, we refer to the metric from RoboBench. As illustrated in the center-left of Figure 9, first, we extract the key steps from the predicted video using Gemini-2.5-flash (Comanici et al., 2025), for instance, 1-grasp(green block), 2-pick up(green block), 3-place(green block, yellow block), and so on, then transform them into a Directed Acyclic Graph(DAG) for further key action extraction and comparison with groundtruth DAG. The detailed metric will be extended in Section 5.3.
@@ 5.4 Overall Benchmark Score
Setup.
For each model $i$ and metric $m$, we map the raw measurement $x_{i,m}$ to a common desirability score $s_{i,m}\in(0,100)$ via a monotone parametric mapping applied after an absolute pre-scaling to $[0,1]$. We then aggregate desirability scores by weighted arithmetic means at both metric-group and overall levels.
Pre-scale to $[0,1]$ with absolute anchors.
Let $L_{m}<U_{m}$ be fixed anchors for metric $m$ (documented per-metric). Define the clipping operator $\mathrm{clip}(u;a,b)=\min\{\max\{u,a\},\,b\}$. We first clamp raw values to $L_{m},U_{m}]$, then linearly map to $[0,1]$; for “higher-is-better” (HIB) metrics:
$\hat{x}_{i,m}^{\mathrm{HIB}}=\frac{\mathrm{clip}(x_{i,m};L_{m},U_{m})-L_{m}}{U_{m}-L_{m}}\in[0,1],$
and for “lower-is-better” (LIB) metrics:
$\hat{x}_{i,m}^{\mathrm{LIB}}=1-\frac{\mathrm{clip}(x_{i,m};L_{m},U_{m})-L_{m}}{U_{m}-L_{m}}\in[0,1].$
We use absolute anchors for two common metrics:
$\textbf{PSNR (HIB):}\quad L_{\text{PSNR}}=0,\ U_{\text{PSNR}}=50\ \ (\text{truncate }x\leq 50\ \text{before scaling});\qquad$
$\textbf{FVD (LIB):}\quad L_{\text{FVD}}=0,\ U_{\text{FVD}}=2000\ \ (\text{truncate }x\leq 2000).$
For other metrics, $(L_{m},U_{m})$ are fixed per protocol (e.g., theoretical bounds for bounded scales, task-specific absolute targets for unbounded ones).
Monotone parametric mappings.
After pre-scaling, we apply a single-parameter monotone transform $f_{m}(\cdot;\theta_{m})$ and then rescale to $(0,100)$:
$s_{i,m}=100\,f_{m}\!\big(\hat{x}_{i,m};\theta_{m}\big),\qquad s_{i,m}\in(0,100).$
We consider the following families (all are strictly increasing on $[0,1]$):
Power (Gamma): $f_{\gamma}(x)=x^{\gamma},\quad\gamma>0;$
Logit temperature: $f_{T}(x)=\sigma\!\big(\mathrm{logit}(x)/T\big),\quad T>0,\ \ \sigma(t)=\tfrac{1}{1+e^{-t}};$
Tanh slope: $f_{\kappa}(x)=\tfrac{1}{2}\big(\tanh(\kappa(2x-1))+1\big),\quad\kappa>0.$
In practice, $\gamma>1$ accentuates the high end, while $T<1$ or $\kappa>1$ expands the mid-range and compresses extremes. For numerical stability with logit we use a small $\varepsilon$ (e.g., $10^{-6}$) and replace $x$ by $\mathrm{clip}(x;\varepsilon,1-\varepsilon)$ only inside $\mathrm{logit}(\cdot)$.
Parameter selection and freezing.
For each metric $m$, $\theta_{m}\in\{\gamma,T,\kappa\}$ is selected on a fixed development set by maximizing the Fisher-$z$ averaged Pearson correlation between $f_{m}(\hat{x};\theta)$ and human ratings across $K$-fold CV; Spearman correlation is used as a tie-breaker. The chosen $\theta_{m}$ is then frozen and applied to all evaluations.
Intra-group averaging (uniform).
Metrics are organized into groups $g$ (e.g., quality, instruction, physical, planning). For model $i$, let $\mathcal{M}_{g}$ be the set of metrics in group $g$ that are available for $i$, and $N_{i,g}=|\mathcal{M}_{g}|>0$. We compute the group score as the simple arithmetic mean:
$G_{i,g}=\frac{1}{N_{i,g}}\sum_{m\in\mathcal{M}_{g}}s_{i,m}.$
Aggregation (weighted arithmetic mean).
Let nonnegative group weights $\{W_{g}\}$ sum to one over the groups available for model $i$. The overall score is
$O_{i}=\sum_{g}\widetilde{W}_{i,g}\,G_{i,g},\qquad\widetilde{W}_{i,g}=\frac{W_{g}}{\sum_{h\in\mathcal{G}_{i}}W_{h}},$
where $\mathcal{G}_{i}=\{g:\,N_{i,g}>0\}$. For an unweighted overall mean across groups, set $W_{g}\equiv 1$.
@@ 6.2 Model Comparsion Experiment
Cogvideo cogvideo 3.29 1.52 1.73 1.30 7.84 38.52 54.09 63.30 2.32 39.56
@@
cosmos1 + Agent cosmos1 35.43 61.07 53.78 8.23 39.63
cosmos2 + Agent cosmos2 49.7 75.96 64.66 11.77 50.53
@@ 6.4.1 Visual Evaluation
We qualitatively evaluate the generalization capabilities of our World Foundation Model across three critical axes: robotic embodiment, task repertoire, and other variations. Cross-embodiment generalization is demonstrated in Figure 14. Proposed world foundation model successfully follows instructions across a diverse set of embodied hardware platforms without any fine-tuning. The model is validated on a wide range of robotic platforms, including the UR5 and Franka industrial manipulators, the Franka Sim simulation, the Agilix parallel-kinematic arm, and the complex Tiangong dexterous hand. This result substantiates our model’s ability to learn an embodiment-agnostic representation of physical interactions, decoupling the task goal from the specific kinematics and dynamics of the embodied platform. Second, Figure 15 illustrates the model’s task generalization, showcasing its proficiency across a repertoire of 15 distinct manipulation skills that range from simple primitives like pull, push, and move, to complex, contact-rich behaviors such as press button, unstack, and tie. The model’s proficiency across this action space highlights its capacity to learn a compositional skill representation rather than merely memorizing individual task solutions. Figure 16 elucidates the model’s remarkable robustness to profound domain and attribute variations. It maintains successful task execution across drastically different visual styles, including photorealistic, pencil sketch, and oil painting. Furthermore, it adeptly handles varied object properties, manipulating both rigid bodies and fluids, from extremely small to regular-sized objects. The model also exhibits invariance to initial state configurations, consistently picking a red pen from a cup regardless of its spatial pose. Collectively, these visualizations provide compelling evidence that the proposed model acquires a truly abstract and foundational understanding of the world, reasoning about underlying physical principles rather than overfitting to superficial contextual cues.
@@ 6.4.2 Quantity Evaluation
Building upon the preceding qualitative analysis, we conduct a rigorous quantitative evaluation across the benchmark’s diverse material properties and physical phenomena. As detailed in Table 4, we assess four models using a suite of five metrics including FVD, WorldScoreBench, PhysGen, DreamSim, and execution quality score. The results unequivocally demonstrate the superiority of Cosmos and WAN over the other models. This performance gap is particularly pronounced in challenging scenarios involving complex dynamics. For instance, Cosmos achieves the highest PhysGen scores, outperforming the next-best model by over 5% on tasks with deformable objects, which suggests a more accurate internal representation of non-rigid body mechanics. In parallel, world foundation model based on Cosmos consistently records the lowest FVD scores in fluid and optical physics simulations, indicating a state-of-the-art capability for rendering visually realistic and temporally coherent complex scenes.